OpenStudy (tootsi123):

A baseball team has home games on Wednesday and Saturday. The two games together earn ​$5000.00 for the team. Wednesday's game generates ​$500.00 less than Saturday's game. How much money was taken in at each​ game?

OpenStudy (tootsi123):

Anyone...

OpenStudy (thereisnousername):

x+y= 5000 x=y-500 substitute the second equation into the first, (y-500)+y=5000 2y-500=5000 2y=5000+500 2y=5500 y=2750 Saturday, x= 2750-500 x=2250 Wednesday, so Saturday was 2750$ and Wednesday was 2250$

OpenStudy (tootsi123):

Thank you! :) So Sat. answer is 2750 and wed is 2250?

OpenStudy (thereisnousername):

yepp :)

OpenStudy (tootsi123):

Thank you! Could u maybe help me with something else?

OpenStudy (tootsi123):

Suppose that a cyclist began a 341 mi ride across a state at the western edge of the​ state, at the same time that a car traveling toward it leaves the eastern end of the state. If the bicycle and car met after 5.5 hr and the car traveled 37.2 mph faster than the​ bicycle, find the average rate of each.